OpenStudy (anonymous):

Aneel has 4 64-ounce cartons of orange juice and 2 48-ounce cans of apple juice. How many ounces of juice does he have altogether?

OpenStudy (whpalmer4):

What is 4*64? What is 2*48? What do you get when you add the two numbers?

OpenStudy (anonymous):

4*64 = 256 , right?

OpenStudy (whpalmer4):

Yes, that's correct. 4*60 = 240, 4*4 = 16, 240+16 = 256

OpenStudy (anonymous):

yes yes yes ( so happy ) but okae i understand the question now, along w/ your help .. THANKS A BUNCH :)

OpenStudy (whpalmer4):

good, glad I could help!